# Ina K. Schieferdecker Technische Universität Berlin, Einsteinufer 25, 10587 Berlin, Germany

### v2

@@ 5 AI-Assisted Model-Driven Software Engineering
    The shift towards the new paradigm of pair modelling in SE, which altogether will turn SE into the next generation of software engineering.
@@
For opportunity (4), the application of AI methods to Big Models is elaborated in first studies as elaborated in table 2 above.
Last but not least for opportunity (5), the pair modelling paradigm naturally evolves from the pair programming paradigm (see e.g. [13]), which has evolved into one of the killer applications of AI in software engineering, where AI-powered tools/agents act as pair programmers and partners in software development (see e.g. [25]).
Based on Big Models and large SE models models for ML, i.e. large models for ML trained with SE models, pair modelling will emerge as a new model-driven software development technique in which a software engineer and an AI tool/agent collaborate in software development. The engineer or tool/agent in the role of the driver writes or improves software artefacts, including models, while the other, in the role of the observer, reviews each element of the software artefact as it is typed into an artefact. The one in the role of observer is also in the role of navigator, considering systemic and strategic aspects of software development: The navigator identifies potential improvements and possible upcoming problems that need to be addressed later on if they are not to be avoided. This allows the driver to focus on the development aspects without losing sight of the cross-cutting and overarching aspects of software development. Like in pair programming in pair modelling, the observer is used as an assurance of the overall resulting software quality and as a guide to high quality software engineering. The two, the driver and observer/navigator, can switch roles. Indeed, the exact interplay of driver and observer/navigator in pair modelling depends on the capabilities of the AI tool and the level of support according to the AI4SE taxonomy, it can provide to the software engineer being in either role.
And as AI is increasingly applied to Big Models, it remains to be seen whether it becomes necessary to include more specifics of modelling in the AI4SE taxonomy, which currently considers modelling as an optional facet in software development, operations and SE processes. At this early stage of development, it is not clear whether modelling needs to be a separate facet or can remain a cross-cutting concern affecting numerous facets within the AI4SE taxonomy.
@@ 6 Related Work
The growing intersection of artificial intelligence (AI) and software engineering (SE) represents a highly active field in research and tool development. The sheer volume of research results has necessitated the creation of surveys [2, 86, 30, 3, 31, 65, 32]that review, sort and structure the state of research in this field. Natural language processing (NLP) techniques, as well as (semi-)formal language processing on large code bases (Big Code) for AI-augmented software engineering, are used to address areas such as code generation, defect detection, code repair, code documentation, software testing, software quality and the challenges related to software security and piracy. Another focus is the application of AI in MDSE [18, 26], examining their use in automating tasks like model generation, transformation, and management, alongside considerations for integration and potential challenges like hallucinations.
It is widely acknowledged that artificial intelligence, and machine learning in particular, is becoming increasingly significant in the field of software engineering. The papers emphasise the potential of AI to automate tasks, enhance code quality, and support developers. It is further noted that AI has the potential to relieve developers from repetitive work, thereby allowing them to focus on higher-level design and problem-solving. Additionally, there is a discussion of the potential of LLMs for code completion, code generation, and other tasks, as well as the challenges and shortcomings of current LLMs. A recurring theme in the literature is the call for more empirical data to support the advancement of this field. Hence, the necessity for a data-driven approach to SE is emphasised by several papers. Also, some papers express concerns about potential biases in AI systems that could have an impact on SE and the resulting software quality.
The papers also exhibit different foci: due to the limited dissemination of MDSE, the majority focusses on the use of AI directly on code, using extensive code bases to train models for code generation and analysis. A smaller portion analyses the application of AI within the context of MDSE, exploring how AI can automate model transformations and facilitate model creation. Furthermore, different taxonomies and classifications are discussed: The papers propose different ways to classify the application of AI in SE, which demonstrates that the field is developing and still lacks a unified framework where AI4SE aims at becoming it. Additionally, there is a divergence of opinion regarding the desired level of automation. Some papers advocate for the utilisation of AI as a tool to assist developers, while others envision the development of fully autonomous systems capable of making decisions without human intervention. Generally, the papers are focused on automating repetitive tasks in software development and reducing errors that may arise from human actions.
More specifically, [32] was one of the first to bridge the gap between AI and SE arguing for their fusion to meet emerging software needs. It explores the potential of integrating AI and SE techniques and tools to create more sophisticated and user-friendly software. [65] discusses how the traditionally separate disciplines of AI and SE are increasingly intertwined. It examines the use of AI techniques to improve SE processes (e.g., project management) and the use of SE methods to develop and manage AI systems (e.g., agent-oriented software engineering). [3] introduces the "naturalness hypothesis" (see section 4) and uses it to justify a data-driven approach to SE. It presents a taxonomy of probabilistic models (code generation, representation, and pattern mining models) and describes their underlying assumptions and applications to SE tasks like code completion, bug detection, and program synthesis. The paper identifies key challenges, such as data sparsity and the need to bridge symbolic programming language representations with ML representations. [86] updates the use of the "naturalness hypothesis" for code generation, completion, translation, refinement, summarisation, defect and clone detection by GitHub Copilot and AlphaCode. It also discusses the challenges and opportunities, including computational cost, quality measurement, software security, piracy concerns, and integration with existing development tools such as Xcode. [2] examines current AI applications across the software development lifecycle, from requirements analysis to maintenance, using tools such as GitHub Copilot and AlphaCode. It analyses the benefits (such as increased productivity and improved code quality) and challenges (such as data quality, ethical concerns and skills gaps) of AI adoption in SE. It also aims to provide a holistic framework for integrating AI responsibly, balancing automation with human oversight, and offers actionable recommendations for practitioners.
Wrt. MDSE, [18] explores the history and future of automation within MDSE. It focuses on the transformative potential of AI for MDSE, addresses challenges such as data scarcity and human-computer interaction, and aims to stimulate further research into the automation capabilities of MDSE through AI and extending its application to different domains and new computing paradigms. [26] examines the burgeoning intersection of LLMs and MDSE, focusing on the automation of tasks such as model repository classification and recommender system development. It explores how LLMs can improve the management, exploration and evolution of entire modelling ecosystems, thereby transforming software engineering practices.
As far as taxonomies for SE and AI are concerned, only two have been made explicit: [31] presents the AI-SEAL taxonomy based on three key facets: point of application (process, product, or runtime), five types of AI technology used (e.g. AI tribes such as symbolist or connectionist), and ten levels of automation (e.g. from "human considers alternatives, makes and implements decision" to "computer makes and implements decision if it feels it should, and informs human only if it feels this is warranted"). In AI4SE, we expand the point/target of the application by detailing the activities involved, we expand the types of AI by adopting a more natural classification to include, for example, agentic AI, and we limit the levels of automation to the more established levels of autonomy when discussing AI applications. In addition, we add the purpose of the AI application, which guides the design, configuration and evaluation of an AI application in SE.
[30] explores the current and future potential of software development bots ("DevBots") as a form of agentic AI. It classifies DevBots according to their purpose, initiation, communication methods, and level of intelligence. It contrasts current DevBots with a vision of an "ideal" DevBot, which would be autonomous, adaptive, and possess both technical and social skills, acting more like artificial team members than mere tools. Like AI4SE, the DevBots taxonomy uses a facet approach to classification, but is restricted to bots, which limits its scope. It also uses a facet for purpose (although this is different from the purpose in AI4SE) and a facet for intelligence (which is limited to only two levels of autonomy/automation). It has two other bot-specific facets for bot initiation and communication.
7 Summary
It is anticipated that the integration of AI will enhance the accessibility of software development to a more heterogeneous user base. By automating tasks, AI can free up developers to focus on higher-level design and problem-solving, leading to a more enjoyable and creative development process. Furthermore, it is regarded as the road towards AI-augmented MDSE bridging the gap between higher-level models and concrete code. However, it is crucial to acknowledge the necessity of addressing the ethical implications of AI in SE, particularly in the context of bias, fairness, transparency, and privacy when developing and deploying AI-augmented software tools. The utilisation of pair programming and/or pair modelling with AI has been identified as a catalyst for enhanced collaborative endeavours between humans and AI. In order to facilitate a more extensive integration of AI into tools, a number of challenges must be surmounted. These include the establishment of interoperability among tools, the creation of standards, and the provision of user-friendly coding and modelling tools. In the future, AI-augmented software lifecycle frameworks will (1) have repetitive and often time-consuming tasks being automated, (2) provide intelligent decision-making by analysing Big Code and/or Big Models for deeper insights, (3) provide enhanced quality of software products by detecting requirements inconsistencies. design flaws, software defects, or other issues, early in the development lifecycle, (4) facilitate management of the increasing complexity of modern systems by providing analysis, monitoring and optimisation tools, and (5) enable continuous improvement by feedback loops and adaptive learning, fostering iterative improvement in software products and software processes.
@@ 7 Summary
Following a review of the current state of model-driven software engineering practice, this paper examines big code on open-source software platforms and the application of AI in software engineering.The novel taxonomy of AI-assisted software engineering, AI4SE, is developed and used to categorise highly cited and recent publications. Furthermore, the concept of Big Models is defined and explored with a view to future opportunities for further adoption of model-driven software engineering, taking advantage of AI gains. Finally, selected research on AI approaches for MDSE is reviewed and categorised using the AI4SE taxonomy.
The novel AI4SE taxonomy aims to classify and understand the different ways in which AI is used in SE, recognising the growing overlap between AI and SE. Future work will focus on the development of Big Models and AI applications in MDSE and/or detail the other activities in SE.

### v3

@@ 5 AI-Assisted Model-Driven Software Engineering
    The shift towards the paradigm of pair modelling in industrial SE practice will altogether turn SE into the next generation of software engineering.
@@
For opportunity (4), Table 2 elaborates the current state of research in more detail.
Finally, regarding opportunity (5), the pair modelling paradigm was initially proposed in [60] when a great deal of work was being conducted on top-down MDSE. In the era of AI-augmented software engineering (SE), the pair modelling paradigm has evolved once again as a natural progression from the pair programming paradigm [21], which has developed into one of AI’s killer applications in SE. Based on big models and "foundation/large models for MDSE" (in analogy to LLM, i.e. large models for ML trained on SE models), AI-supported pair modelling may also become a central AI application for MDSE: AI-powered tools and agents could act as pair modellers and partners in software development. See, for example, [34] for similarities with pair programming. The engineer or tool/agent in the role of the driver writes or improves software artefacts, including models, while the other, in the role of the observer, reviews each element of the software artefact as it is typed into an artefact. The one in the role of observer is also in the role of navigator, considering systemic and strategic aspects of software development: The navigator identifies potential improvements and possible upcoming problems that need to be addressed later on if they are not to be avoided. This allows the driver to focus on the development aspects without losing sight of the cross-cutting and overarching aspects of software development. As with pair programming, pair modelling uses an observer to ensure the quality of the final software and guide high-quality software engineering. The driver and observer/navigator can switch roles. Indeed, the exact interplay of driver and observer/navigator in pair modelling depends on the capabilities of the AI tool and the level of automation support, it can provide to the software engineer being in either role.
@@ 6 Related Work
As this paper’s main contribution is the ai4se taxonomy, which was developed through a literature review of AI for SE (including MDSE), the related work section reviews literature survey approaches and taxonomy developments in the emerging field of AI applications in SE. Classifying research papers within this taxonomy also led to a discussion of opportunities in AI for MDSE. Therefore, the related work section also covers roadmaps on this topic.
Systematic literature reviews are used to explore the status of a research field by examining the results of published research presented in peer-reviewed papers that are available online. Numerous SLRs on AI for SE have been published. While some categorise the reviewed research, none have developed a taxonomy for classifying the field of AI for SE. Table 3 presents a selection of recent SLRs on AI for SE.
Paper Summary Difference to ai4se SLR
[5] This paper focuses on the benefits of AI for SE, such as increased productivity, improved code quality and faster development, and on key impact areas such as automated coding, intelligent debugging and predictive maintenance. The paper supports its findings with case studies and theoretical models from various activities in SE. SLR with focus on case studies.
[39] This study provides a quantitative assessment of the impact of AI in SE phases. It suggests that AI can improve the accuracy of planning and requirement engineering, as well as enhancing the efficiency of the latter and the software design. Follow-up quantitative analysis of a SLR.
[57] This study analyses LLMs for SE by characterising their features, reviewing data curation and training processes, and evaluating their performance and success in application to SE. SLR on LLMs for SE.
[81] This study examines the main challenges of applying AI to project management, particularly during the planning stage. SLR on AI for project planning.
[123] This paper provides an overview of the general workflow of learning-based automated program repair techniques. It outlines related empirical studies and provides evaluation metrics. SLR on AI for program repair.
[38] This paper aims at identifying the AI techniques most often used in software development and their effects on the accuracy and efficiency of software development activities. SLR on software development.
[117] This paper investigates the software testing tasks for which LLMs are typically employed, the most prevalent LLMs, the types of prompt engineering employed alongside these LLMs, and the associated techniques. SLR on LLMs for testing.
[66] This study analysis the advances of AI for SE, the potentials for future development as well as the risks of AI application to SE. SLR combined with qualitative expert interviews.
[120] This paper provides a review of transformer-based LLMs that have been trained using big code and are used for NLP techniques in AI-assisted programming. SLR on NLP utilization for programming.
[118] This study examines the complexities involved in applying ML/DL solutions to SE. It also explores how these issues affect the reproducibility and replicability of ML/DL applications in SE. SLR on ML/DL for SE.
[82] This study explores the extent to which explainable AI has been investigated within the SE community. Software maintenance, and defect prediction in particular, is the area of SE that has received the most attention in terms of the stages and tasks being studied. SLR on explainable AI for SE.
[42] This paper reviews the use of LLMs in SE and highlights the important role of a hybrid approach combining traditional SE and LLMs in achieving reliable, efficient and effective LLM-based SE. SLR on LLMs for SE.
Table 3: SLRs on AI for SE (sorted by publication year)
Another important area of research is the development of roadmaps. Like SLRs, these seek to evaluate and categorise developments thus far. In addition, however, they project these developments into the future. The roadmaps presented in Table 4 were recently published.
Paper Summary Difference to ai4se taxonomy
[109] This paper presents a vision for an AI-driven software development framework. The main actors in this framework are software engineers, including developers, architects and testers, as well as a generic AI system, such as an LLM. It discusses the challenges for requirements engineering, software testing, development and testing, and maintenance, which is more coarse-grained than ai4se.
[1] This study examines how AI is reshaping the role of humans within the software ecosystem. It explores the challenges and opportunities arising from the interaction between technical and human factors in human-centred workflows. This categorization of AI for SE focuses on how AI can enhance SE, render technologies and approaches obsolete, overturn common practice, retrieve past results, and change workflows. Human factors have yet to be incorporated into ai4se, and could be considered a new dimension.
[4] This study highlights three key challenges for AI in SE: the use of GenAI and LLMs for engineering large software systems; the requirement for large, unbiased datasets and benchmarks for training and evaluating DL and LLMs for SE; and the need for a new code of digital ethics for applying AI in SE. By analysing the challenges in classical and upcoming SE processes, the relevance of human factors in SE is also emphasised (see also [1]).
[35] This paper presents an overview of current LLM applications in MD[S]E and a roadmap for the deployment of LLMs to enhance the management, exploration, and evolution of modelling ecosystems. The status of LLMs for MDSE is presented as a feature model. In particular, the more detailed modelling tasks could be considered new facets of MDSE in ai4se.
[77] This paper explores the potential applications of GenAI in model-driven software maintenance and evolution. As represented in ai4se, this paper explores how MDSE can be advanced through AI, particularly GenAI. It emphasises that AI can assist modellers, enhance their capabilities, and facilitate reasoning and automation in MDSE.
[16] This study analyses the infusion of AI techniques into DevOps and software security-related tasks. It also presents a call for action towards GenAI-based agents for SE. The study emphasises the importance of human factors, security and trust considerations, all of which are candidate extensions for ai4se.
[72] This paper reviews the development stages of AI for SE, highlighting trust and synergy as two key challenges that must be overcome before AI for SE tools can act as intelligent, responsible workmates. The paper proposes research into trust-aware, privacy-aware, licence-aware, attack-resistant, secure and workflow-aware AI for SE solutions. This could potentially lead to new research being classified in the ai4se taxonomy.
[85] With the help of an SLR, this paper identifies 78 open research questions across 11 SE areas for a GenAI-based SE solution. A research agenda is derived that highlights the need for solutions in the areas of dependability and accuracy, data accessibility and transparency, and sustainability aspects of GenAI solutions for AI. These requirements could potentially lead to new research being classified in the ai4se taxonomy. Moreover, the ’quality’ aspects of AI for SE solutions, such as accuracy, robustness, transparency and sustainability, have the potential to be an extension of ai4se.
Table 4: Research roadmaps on AI for SE (sorted by publication year)
Last but not least, some papers have investigated the development of a taxonomy for AI for SE, often focusing on a subfield of SE or the AI techniques used, as described in Table 5.
Paper Summary Difference to ai4se taxonomy
[111] The types of interaction between software developers and AI are the focus of this taxonomy. Human factors are a candidate extension for ai4se.
[80] The DAnTE taxonomy focusses on the degree of automation of SE tools in general and defines six levels, including no automation, informer, suggester, and local, global and full generators. It also reviews AI for SE in particular. DAnTE considers three types of generator, whereas ai4se differentiates between partial and full generators only. The differences between ’global’ and ’full’ are subtle, and no tool currently addresses these levels.
[124] This taxonomy organizes 57 most frequently used NLP techniques in requirements engineering (RE). This NLP classification could be used to further refine the AI types in ai4se.
[101] This systematic mapping study examines the use of four types of AI (and their combinations) in the development, deployment and maintenance of software. While the AI types and the SE activities in ai4se are more detailed, this mapping’s approach is similar. Still, for ai4se, it could be considered to make the (AI) training a separate activity.
[74] This systematic mapping study examines the use of AI in component-based SE and its potential to support this field. Not only does it map AI approaches to SE activities, it also identifies the problems being addressed. ai4se considers SE in general rather than component-based SE specifically. Nevertheless, some of the AI types and targets, which are the ’problems’ in this study, may be considered for inclusion.
[41] This study focuses on bots that support software development (DevBots). It classifies contemporary DevBots using a facet-based taxonomy and outlines the ideal characteristics of future DevBots. The main facets are purpose, initiation, communication and intelligence. While the main facets ’purpose’ and ’intelligence’ essentially correspond to the ai4se dimensions ’purpose’ and ’AI type’, ’initiation’ and ’communication’ could be considered for inclusion in ai4se as part of the human factors.
[43] This study introduces the AI-SEAL taxonomy, which classifies AI solutions for SE based on their application point, the type of AI technology employed, and the degree of automation they offer. AI-SEAL uses three high-level SE processes, five AI tribes for categorising AI types and ten automation levels to determine the level of automation support. In contrast, ai4se provides a detailed view of SE, categorising AI types into seven categories and automation levels into five.
Table 5: Taxonomies on AI for SE (sorted by publication year)
The SLRs, roadmaps, and taxonomies presented in Tables 3, 4, and 5 show that the intersection between AI and SE is a rapidly growing area of research and tool development. AI support in SE is used to address areas such as code generation, defect detection, code repair, code documentation, software testing, and challenges relating to software quality, security, and piracy. The application of AI in MDSE to rail guard the application of AI by models is another focus.
Only a few taxonomies for SE and AI have been made explicit so far. They address a specific subfield of SE only or focus on classifying the AI methods and techniques used. Still, all of these taxonomies essentially align with ai4se. In addition, the taxonomies in [43, 80] and [41] highlight the need to detail human factors when analysing the potentials of AI for SE. In future work, ai4se is to be extended accordingly.
7 Summary and Outlook
@@
Following a review of the current state of model-driven software engineering practice, this paper examines big code on open-source software platforms and the application of AI in software engineering. The novel ai4se taxonomy is developed and employed to categorise recent and highly cited publications. This taxonomy aims to classify and understand the various ways in which AI is employed in software engineering, acknowledging the increasing overlap between the two fields. Furthermore, the concept of big models is defined and explored to identify opportunities for further adoption of model-driven software engineering that capitalises on AI progress. Finally, selected research on AI approaches for MDSE is reviewed and categorised using the ai4se taxonomy. Future work will focus on contributing to the development of big models and AI applications in MDSE, as well as detailing other SE activities within the taxonomy.
With regard to the research questions provided for the SLR, which formed the basis for the development of the ai4se taxonomy, the answers are as follows: The SLR revealed current research on AI for SE and for MDSE as shown in Table 1 and Table 2. Software development and model-driven software design are the primary focus of AI support in SE, but also other SE and MDSE activities are being studied. The ai4se taxonomy can classify the identified unique research, that address all dimensions and their facets, but not general purpose AI (T5) or AI-driven full automation (L4). Subsymbolic AI (T2) and generative AI (T3) are often used, agentic AI (T4) is on a rise. The automation levels reached are most often AI-assisted option (L1) and AI-assisted selection (L2).
Overall, integrating AI into SE is expected to make the field more accessible to a more diverse range of users. Increasing the automation of SE tasks frees up developers to focus on higher-level design and problem solving, thereby making the development process more enjoyable and creative. Furthermore, AI-augmented MDSE is regarded as the way forward, seamlessly bridging the gap between higher-level models and concrete code. Large models will facilitate the model-based training of GenAI and Agentic AI across the entire MDSE landscape.
However, it is also crucial to acknowledge the necessity of addressing the ethical implications of AI in SE, particularly with regard to bias, fairness, transparency and privacy, when developing and deploying AI-augmented software tools. Using pair programming and pair modelling with AI has been identified as a way to enhance collaboration between humans and AI, and to address ethical considerations.
Moreover, in order to allow for greater integration of AI into tools, a number of challenges must be overcome. These include establishing interoperability between tools, creating standards and providing user-friendly coding and modelling tools. In the future, AI-augmented software lifecycle frameworks will
  1. 1.
    automate repetitive and often time-consuming tasks,
  2. 2.
    provide intelligent decision-making by analysing big code and big models for deeper insights,
  3. 3.
    enhance the quality of software products by detecting inconsistencies in requirements, design flaws and software defects early in the development lifecycle,
  4. 4.
    facilitate management of the increasing complexity of modern systems by providing analysis, monitoring and optimisation tools, and
  5. 5.
    enable continuous improvement through feedback loops and adaptive learning, fostering iterative improvement in software products and processes.
In order for the community to follow the further evolution of AI for SE and MDSE, the ontology and paper classification along the ai4se taxonomy have been made available to the public under a CC-BY-SA licence for use and extension as required.